OpenStudy (anonymous):

For the inverse variation equation of xy = 19, complete the statement, y________as x decreases. increases decreases remains constant I think it remains constant but I also think it increases, im not sure please explain

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (goformit100):

decreases 

jimthompson5910 (jim_thompson5910):

let's say x = 1 xy = 19 1*y = 19 y = 19 now let's say x increases to 19 xy = 19 19*y = 19 y = 19/19 y = 1 so as x increased, did y increase, decrease, or stay the same?

OpenStudy (anonymous):

stays the same?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x goes from 1 to 19 y goes from 19 to 1 see this?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so as x increases, y decreases if x decreases, then y increases (everything flips)

jimthompson5910 (jim_thompson5910):

so the final answer is "increases"

jimthompson5910 (jim_thompson5910):

remember with inverse, as one thing goes up, the other goes down (or vice versa)

OpenStudy (anonymous):

okay so it will always be one or another not "stay the same"

jimthompson5910 (jim_thompson5910):

correct, nothing will stay the same if one variable changes

OpenStudy (anonymous):

okay thank you!

jimthompson5910 (jim_thompson5910):

yw